Mitochondria are the powerhouses of eukaryotic cells. Roughly, an average size mitochondrion can be modeled as a cylinder of radius 0.57 μm and length 2.2 μm. What is the total surface area of the mitochondrion, including the two caps? 010 μη? 9.9 μη?
Answer #1

S = 2*\pi*R(L+R) =2*\pi*.57(2.2+.57) = 9.92 \mum2